# Witnessing superpositions of causal orders by weak measurements at a given time

### v2

@@ Abstract
The questions we raise in this letter are as follows: What is the most general representation of a quantum state at a single point in time? Can we adapt the current formalisms to situations where the order of quantum operations is coherently or incoherently superposed? If so, what are the relations between the state at a given time and the uncertainty in the order of events before and after it? Establishing the relationship between two-state vector formalism and pseudo-density operators, we introduce the notion of a single-time pseudo-state. The tomographic construction of single-time pseudo-states is possible by ideal or weak measurements. We demonstrate that the eigenspectrum obtained from weak measurements enables us to discriminate between some coherent and incoherent superpositions of causal orders in which the same pre- and post-selection events interchange with a non-zero probability. Finally, we discuss some possible experimental realizations in existing photonic setups.
@@
In 1995, Reznik and Aharonov expressed the two-state in the form of a direct product of $|\psi\rangle$ and $\langle\phi|$ [3], aiming to formulate the statistics of ideal and weak measurements [5, 6] in a unified manner. However, $|\psi\rangle\langle\phi|$ is generally not a Hermitian, positive-semidefinite, or normalized operator. In a recent reformulation, Silva et al. represented (2) as ${}_{t_{2}}\!\langle\phi|\otimes|\psi\rangle_{t_{1}}$ [7]. To generalize the concept of two-state for incoherent mixtures $\{p^{r},\Psi^{r}\}$, they defined a density vector such that $\eta=\sum_{r}p^{r}\Psi^{r}\otimes\Psi^{r\dagger}$, where $\Psi^{r}=\sum_{i,j}\alpha^{r}_{ij}\,{}_{t_{2}}\!\langle i|\otimes|j\rangle_{t_{1}}$ is a coherent superposition of two-states. This reformulation facilitated the demonstration that a subset of pre- and post-selected states corresponds to process matrices [9], which are conventional mathematical objects used to characterize indefinite causal orders [10, 9, 11, 12, 13, 14, 15] in the literature. Note that the term indefinite causal structure encompasses not only the superposition of different causal orders but also encompasses the superposition of various causal relations, including cause-effect and common cause relations [16].
Our objective is to employ quantum state tomography to identify and characterize alternative mathematical objects that can encode measurement statistics for pre- and post-selected systems. By doing so, we aim to maintain a connection with the density matrix formalism as much as possible while describing these systems. One of us recently developed a similar approach in Ref. [17] that treats space and time equally in quantum mechanics. This approach assumes that the state of time-like separated systems can be reconstructed by a complete multi-time tomography. Moreover, the presence of negative eigenvalues in the resultant multi-time pseudo-density matrix/operator (PDO) is interpreted as evidence of temporal quantum correlations shared between the systems [17, 18, 19]. Recently, mappings between PDOs, two-states, and process matrices have also been identified [20]. Here, we depart from these previous studies by considering complete single-time tomography of the system. Furthermore, we extend the tomographic construction to incorporate weak measurements [5, 6]. The resultant single-time PDO enables us to discriminate between some coherent and incoherent superpositions of causal orders in which the identical pre- and post-selection events interchange with a non-zero probability.
@@
If the intermediate measurement were ideal like pre- and post-selection measurements, even fixed causal order would result in a pseudo-state with a negative eigenvalue as shown in Appendix A. However, any mixed causal order cannot be distinguishable from fixed causal orders in this way. On the other hand, the negativity of the eigenvalue of $\Gamma_{\upharpoonleft\!\downharpoonright}(t)$ can be used as an experimental witness of incoherent mixture of causal orders unless $|\psi\rangle$ and $|\phi\rangle$ are eigenvectors of two different Pauli operators (see Appendix B).
@@
The coherent superposition of two opposite causal orders is referred to the indefinite causal order in the literature [10, 9, 11, 12, 14, 16, 15]. Such superpositions can be generated by a quantum supermap called quantum SWITCH [10, 11], which requires the presence of an ancillary two-level quantum system that acts as a control qubit. When the initial state of control qubit is changed from $|0\rangle_{c}$ to $|1\rangle_{c}$, the system of interest experiences two subsequent physical processes in a reversed order. Therefore, if the control qubit is prepared in a superposition state $(|0\rangle_{c}+|1\rangle_{c})/\sqrt{2}$, then a coherent superposition of the opposite causal orders can be created.
@@
An optical implementation of quantum SWITCH was recently reported in Ref. [13]. In this context, the photonic setups we proposed above can be extended to include quantum SWITCH.
Alternatively, an entanglement swapping protocol, as proposed by Aharonov and coworkers [2, 4] can also be utilized to generate a coherent superposition of opposite causal orders. This protocol similarly necessitates the presence of an ancillary two-level quantum system, akin to the SWITCH. Initially, the system and the ancilla are prepared in the entangled state $(|\Psi\rangle=|\psi\rangle|0\rangle_{c}+|\phi\rangle|1\rangle_{c})/\sqrt{2}$ at time $t_{1}$. Subsequently, the ancilla remains shielded from any interaction until the post-selection measurement at time $t_{2}$, which projects the joint state onto $(|\Phi\rangle=|\phi\rangle|0\rangle_{c}+|\psi\rangle|1\rangle_{c})/\sqrt{2}$. This projection effectively swaps the entanglement between the system and the ancilla to the two-state of the system starting from time $t_{2}$ onward.
Thermal witnessing of indefinite causal orders.— Let us assume that the Hamiltonian of the system of interest is $H_{S}=E_{0}|0\rangle\langle 0|+E_{1}|1\rangle\langle 1|$. Consider the following pre- and post-selected states which are purified by including some ancillary degrees of freedom $A_{j}$
$|\Psi(t_{j})\rangle$ $=|\omega_{j}\rangle\equiv\sqrt{p_{0}}|0\rangle_{S}|0\rangle_{A_{j}}+i^{(j-1)}\sqrt{p_{1}}|1\rangle_{S}|1\rangle_{A_{j}},\,$ (17)
@@
Consider the construction of a pseudo-state characterizing the system at an intermediate time instant $t$ between $t_{1}$ and $t_{1}$ through weak measurement-based tomography. Can the effective temperature of this single-time pseudo-state offer insights into the order of operations carried out at $t_{1}$ and $t_{2}$? Indeed, $\Gamma_{\uparrow}(t)=\text{tr}_{A_{1}A_{2}}[|\omega_{1}\rangle\langle\omega_{2}|/\langle\omega_{2}|\omega_{1}\rangle]$ and $\Gamma_{\downarrow}(t)=\text{tr}_{A_{1}A_{2}}[|\omega_{2}\rangle\langle\omega_{1}|/\langle\omega_{1}|\omega_{2}\rangle]$ represent effective Gibbs states characterized by complex inverse temperatures. Neglecting unitary time evolution between $t_{1}$ and $t_{2}$, we derive $\beta_{\uparrow}(t)=\beta+i\,\pi/(2\,\Delta E)$ and $\beta_{\downarrow}(t)=\beta-i\,\pi/(2\,\Delta E)$ where $\Delta E=E_{1}-E_{0}$. Thus, the imaginary component of the effective temperature at $t$ can discern between two opposite causal orders, regardless of the specific values $\beta$ at $t_{1}$ and $t_{2}$ (see Fig. 3-a and -b).
@@
Open problems.— The conventional formalism employed in the literature to explore the quantum superposition of causal orders is grounded in process matrices [9]. Two-time two-states and multi-time pseudo-states, as they can be mapped onto process matrices, are also capable of modeling indefinite causal orders [8, 20]. In this letter, we focused on weak measurements that can be performed within different causal structures at a single time. We demonstrated that single-time pseudo-states, which only describe the statistics of these measurements, can also provide information about the causal structure. Given their relatively compact size compared to multi-time two-states and multi-time pseudo-states, these pseudo-states hold promise in effectively modeling more intricate causal structures.

### v3

# Witnessing superpositions of causal orders before the process is completed

@@ Abstract
The questions we raise in this letter are as follows: What is the most general representation of a quantum state at a single point in time? Can we adapt the current formalisms to situations where the order of quantum operations is coherently or incoherently superposed? If so, what are the relations between the state at a given time and the uncertainty in the order of events before and after it? Establishing the relationship between two-state vector formalism and pseudo-density operators, we introduce the notion of a single-time pseudo-state. The tomographic construction of single-time pseudo-states is possible by ideal or weak measurements. We demonstrate that the eigenspectrum obtained from weak measurements enables us to discriminate between some coherent and incoherent superpositions of causal orders in pre- and post-selected systems before the process is completed. Finally, we discuss some possible experimental realizations in existing photonic setups.
@@
In 1995, Reznik and Aharonov expressed the two-state in the form of a direct product of $|\psi\rangle$ and $\langle\phi|$ [3], aiming to formulate the statistics of ideal and weak measurements [5, 6] in a unified manner. However, $|\psi\rangle\langle\phi|$ is generally not a Hermitian, positive-semidefinite, or normalized operator. In a recent reformulation, Silva et al. represented (2) as ${}_{t_{2}}\!\langle\phi|\otimes|\psi\rangle_{t_{1}}$ [7]. To generalize the concept of two-state for incoherent mixtures $\{p^{r},\Psi^{r}\}$, they defined a density vector such that $\eta=\sum_{r}p^{r}\Psi^{r}\otimes\Psi^{r\dagger}$, where $\Psi^{r}=\sum_{i,j}\alpha^{r}_{ij}\,{}_{t_{2}}\!\langle i|\otimes|j\rangle_{t_{1}}$ is a coherent superposition of two-states. This reformulation facilitated the demonstration that a subset of pre- and post-selected states corresponds to process matrices [9], which are conventional mathematical objects used to characterize indefinite causal orders (ICOs) in the literature.
ICOs were initially proposed based on the notion that quantum gravity should be a probabilistic theory with a dynamic causal structure [10, 11, 12]. Hardy also suggested that computers operating without a definite causal structure could surpass conventional computers in performance [13]. The introduction of the quantum SWITCH has theoretically demonstrated in Refs. [14, 15] that such indefiniteness indeed offers computational advantages. The resource value of ICOs  [^1] Note that the term indefinite causal structure encompasses not only the superposition of different causal orders but also encompasses the superposition of various causal relations, including cause-effect and common cause relations [16]. has since been shown in various fields, including quantum computation [17, 18, 19, 20, 21], quantum communication [22, 23, 24], quantum metrology [25, 26], and quantum thermodynamics [27, 28, 29]. However, to efficiently harness this resource, it must first be identified. Current ICO witnesses in the literature [30, 31, 32] can only detect this resource after the processes have been completed.
Our objective is to employ quantum state tomography to identify and characterize alternative mathematical objects that can encode measurement statistics for pre- and post-selected systems. By doing so, we aim to detect ICOs in these systems before the process is completed. One of us recently developed a similar approach in Ref. [33] that treats space and time equally in quantum mechanics. This approach assumes that the state of time-like separated systems can be reconstructed by a complete multi-time tomography. Moreover, the presence of negative eigenvalues in the resultant multi-time pseudo-density matrix/operator (PDO) is interpreted as evidence of temporal quantum correlations shared between the systems [33, 34, 35]. Recently, mappings between PDOs, two-states, and process matrices have also been identified [36]. Here, we depart from these previous studies by considering complete single-time tomography of the system. Furthermore, we extend the tomographic construction to incorporate weak measurements [5, 6]. The resultant single-time PDO enables us to discriminate between some coherent and incoherent superpositions of causal orders in which the identical pre- and post-selection events interchange with a non-zero probability.
@@
If the intermediate measurement were ideal like pre- and post-selection measurements, even fixed causal order would result in a pseudo-state with a negative eigenvalue as shown in Appendix A. However, any mixed causal order cannot be distinguishable from fixed causal orders in this way.
The coherent superposition of causal orders can be experimentally realized in photonic setups using a quantum SWITCH [31, 32]. In these setups, if the initial state of the control qubit could be considered mixed, an incoherent mixture of causal orders could also be experimentally created. However, since the control qubit in these setups is typically represented by the path [31] or polarization [32] degree of freedom of a photon, preparing an initial mixed state can be complicated. Therefore, we propose alternative experimental setups for creating an incoherent mixture of causal orders.
@@
The coherent superposition of two opposite causal orders can be generated by a quantum supermap called quantum SWITCH [14, 15], which requires the presence of an ancillary two-level quantum system that acts as a control qubit. When the initial state of control qubit is changed from $|0\rangle_{c}$ to $|1\rangle_{c}$, the system of interest experiences two subsequent physical processes in a reversed order. Therefore, if the control qubit is prepared in a superposition state $(|0\rangle_{c}+|1\rangle_{c})/\sqrt{2}$, then a coherent superposition of the opposite causal orders can be created.
@@
Different optical implementations of quantum SWITCH were recently reported in Refs. [31, 32]. In this context, the photonic setups we proposed above can be extended to include quantum SWITCH. Alternatively, an entanglement swapping protocol, as proposed by Aharonov and coworkers [2, 4] can also be utilized to generate a coherent superposition of opposite causal orders. This protocol similarly necessitates the presence of an ancillary two-level quantum system, akin to the SWITCH. Initially, the system and the ancilla are prepared in the entangled state $(|\Psi\rangle=|\psi\rangle|0\rangle_{c}+|\phi\rangle|1\rangle_{c})/\sqrt{2}$ at time $t_{1}$. Subsequently, the ancilla remains shielded from any interaction until the post-selection measurement at time $t_{2}$, which projects the joint state onto $(|\Phi\rangle=|\phi\rangle|0\rangle_{c}+|\psi\rangle|1\rangle_{c})/\sqrt{2}$. This projection effectively swaps the entanglement between the system and the ancilla to the two-state of the system starting from time $t_{2}$ onward.
Thermal witnessing of indefinite causal orders.— Let us assume that the Hamiltonian of the system of interest is $H_{S}=E_{0}|0\rangle\langle 0|+E_{1}|1\rangle\langle 1|$. Consider the following pre- and post-selected states which are purified by including some ancillary degrees of freedom $A$
$|\Psi(t_{j})\rangle$ $=|\omega_{j}\rangle\equiv\sqrt{p_{0}}|0\rangle_{S}|0\rangle_{A}+i^{(j-1)}\sqrt{p_{1}}|1\rangle_{S}|1\rangle_{A},\,$ (17)
@@
Consider the construction of a pseudo-state characterizing the system at an intermediate time instant $t$ between $t_{1}$ and $t_{2}$ through weak measurement-based tomography. Can the effective temperature of this single-time pseudo-state offer insights into the order of operations carried out at $t_{1}$ and $t_{2}$? Indeed, $\Gamma_{\uparrow}(t)=\text{tr}_{A_{1}A_{2}}[|\omega_{1}\rangle\langle\omega_{2}|/\langle\omega_{2}|\omega_{1}\rangle]$ and $\Gamma_{\downarrow}(t)=\text{tr}_{A_{1}A_{2}}[|\omega_{2}\rangle\langle\omega_{1}|/\langle\omega_{1}|\omega_{2}\rangle]$ represent effective Gibbs states characterized by complex inverse temperatures. Neglecting unitary time evolution between $t_{1}$ and $t_{2}$, we derive $\beta_{\uparrow}(t)=\beta+i\,\pi/(2\,\Delta E)$ and $\beta_{\downarrow}(t)=\beta-i\,\pi/(2\,\Delta E)$ where $\Delta E=E_{1}-E_{0}$. Thus, the imaginary component of the effective temperature at $t$ can discern between two opposite causal orders, regardless of the specific values $\beta$ at $t_{1}$ and $t_{2}$ (see Fig. 3-a and -b).
@@
Open problems.— The conventional formalism employed in the literature to explore the quantum superposition of causal orders is grounded in process matrices [9]. Two-time two-states and multi-time pseudo-states, as they can be mapped onto process matrices, are also capable of modeling indefinite causal orders [8, 36]. In this letter, we focused on weak measurements that can be performed within different causal structures at a single time. We demonstrated that single-time pseudo-states, which only describe the statistics of these measurements, can also provide information about the causal structure. Given their relatively compact size compared to multi-time two-states and multi-time pseudo-states, as well as their ability to identify the causal order prior to the completion of the process, these pseudo-states hold promise in effectively modeling more intricate causal structures.
@@
FIG. 3: Thermal witnessing of indefinite causal orders. a) When the control qubit (C) is in the $|0\rangle$ state, the system of interest (S) and an ancillary system (A) are projected onto the entangled states of $|\omega_{1}\rangle$ ($|\omega_{2}\rangle$) at time $t_{1}$ ($t_{2}$). Through these projections, S acquires an effective temperature $T_{S}$. If a weak measurement-based tomography is conducted at an intermediate time $t$ to construct the pseudo-state of S, the effective temperature gains an imaginary component. b) When C is in the $|1\rangle$ state, the order of the projection operations at times $t_{1}$ and $t_{2}$ changes. This change alters the imaginary component of the effective temperature at time $t$. c) In the case where C exists in $|0\rangle$ and $|1\rangle$ states with equal probabilities, an incoherent mixture of fixed causal orders in (a) and (b) is obtained. Consequently, the effective temperature at time $t$ reduces to half of the temperature at times times $t_{1}$ and $t_{2}$ changes. d) If C exists in a superposition of $|0\rangle$ and $|1\rangle$ states, it results in a quantum superposition of fixed causal orders in (a) and (b). Notably, in this scenario, the effective temperature at time $t$ remains unaffected by the temperature observed at times $t_{1}$ and $t_{2}$, persistently holding at 00. In this example, the emergence of fixed causal orders with complex-valued temperatures and incoherent/coherent superpositions with real-valued temperatures can be attributed to the methodology utilized in weak measurement-based tomography. Notably, this approach focuses exclusively on a subsystem, S, rather than the entire system. Expanding tomography to encompass the entangled SA system reveals a distinct pattern: determinants of fixed causal order pseudo-states register as zero, while others exhibit negative determinants. This observation underscores the significance of considering the interplay between causal and spatial correlations.